Stocks A and B have the following probability distributions of expected future returns:

Probability     A     B
0.1 (8 %) (40 %)
0.2 4 0
0.4 14 19
0.2 18 29
0.1 38 45
  1. Calculate the expected rate of return, , for Stock B ( = 13.00%.) Do not round intermediate calculations. Round your answer to two decimal places.

      %

  2. Calculate the standard deviation of expected returns, σA, for Stock A (σB = 21.95%.) Do not round intermediate calculations. Round your answer to two decimal places.

      %

    Now calculate the coefficient of variation for Stock B. Do not round intermediate calculations. Round your answer to two decimal places.

    Is it possible that most investors might regard Stock B as being less risky than Stock A?

    1. If Stock B is more highly correlated with the market than A, then it might have a lower beta than Stock A, and hence be less risky in a portfolio sense.
    2. If Stock B is more highly correlated with the market than A, then it might have the same beta as Stock A, and hence be just as risky in a portfolio sense.
    3. If Stock B is less highly correlated with the market than A, then it might have a lower beta than Stock A, and hence be less risky in a portfolio sense.
    4. If Stock B is less highly correlated with the market than A, then it might have a higher beta than Stock A, and hence be more risky in a portfolio sense.
    5. If Stock B is more highly correlated with the market than A, then it might have a higher beta than Stock A, and hence be less risky in a portfolio sense.

    -Select-IIIIIIIVVItem 4

  3. Assume the risk-free rate is 3.5%. What are the Sharpe ratios for Stocks A and B? Do not round intermediate calculations. Round your answers to four decimal places.

    Stock A:

    Stock B:

    Are these calculations consistent with the information obtained from the coefficient of variation calculations in Part b?

    1. In a stand-alone risk sense A is less risky than B. If Stock B is less highly correlated with the market than A, then it might have a higher beta than Stock A, and hence be more risky in a portfolio sense.
    2. In a stand-alone risk sense A is more risky than B. If Stock B is less highly correlated with the market than A, then it might have a lower beta than Stock A, and hence be less risky in a portfolio sense.
    3. In a stand-alone risk sense A is more risky than B. If Stock B is less highly correlated with the market than A, then it might have a higher beta than Stock A, and hence be more risky in a portfolio sense.
    4. In a stand-alone risk sense A is less risky than B. If Stock B is more highly correlated with the market than A, then it might have the same beta as Stock A, and hence be just as risky in a portfolio sense.
    5. In a stand-alone risk sense A is less risky than B. If Stock B is less highly correlated with the market than A, then it might have a lower beta than Stock A, and hence be less risky in a portfolio sense.
Answer #1

1.
=0.1*(-40%)+0.2*0%+0.4*19%+0.2*29%+0.1*45%=13.90%

2.
=SQRT(0.1*(-8%-13%)^2+0.2*(4%-13%)^2+0.4*(14%-13%)^2+0.2*(18%-13%)^2+0.1*(38%-13%)^2)=11.3225438837745%

3.
=21.95%/13.90%
=1.57914

4.
If Stock B is less highly correlated with the market than A, then it might have a lower beta than Stock A, and hence be less risky in a portfolio sense.

5.
=(13%-3.5%)/11.3225438837745%
=0.83903

6.
(13.90%-3.5%)/21.95%
=0.47380

7.
In a stand-alone risk sense A is more risky than B. If Stock B is less highly correlated with the market than A, then it might have a lower beta than Stock A, and hence be less risky in a portfolio sense.